# Efficient Quantum Pseudorandomness from Hamiltonian Phase States

### v2

@@ Abstract
In this work, we seek to make progress on both of these fronts simultaneously—by decoupling quantum pseudorandomness from classical cryptography altogether. We introduce a quantum hardness assumption called the Hamiltonian Phase State ($\mathsf{HPS}$) problem, which is the task of decoding output states of a random instantaneous quantum polynomial-time (IQP) circuit. Hamiltonian phase states can be generated very efficiently using only Hadamard gates, single-qubit $Z$ rotations and CNOT circuits. We show that the hardness of our problem reduces to a worst-case version of the problem, and we provide evidence that our assumption is plausibly fully quantum; meaning, it cannot be used to construct one-way functions. We also show information-theoretic hardness when only few copies of $\mathsf{HPS}$ are available by proving an approximate $t$-design property of our ensemble. Finally, we show that our $\mathsf{HPS}$ assumption and its variants allow us to efficiently construct many pseudorandom quantum primitives, ranging from pseudorandom states, to quantum pseudoentanglement, to pseudorandom unitaries, and even primitives such as public-key encryption with quantum keys. Along the way, we analyze a natural iterative construction of pseudorandom unitaries which resembles a candidate of Ji, Liu, and Song (CRYPTO’18).
@@ 2.2 Overview of our Results
  • (Evidence of worst-case hardness:) The learnability of ensembles of phase states has been studied extensively in recent work [arunachalam2023optimal], and has been found to have exponential time complexity in the worst case (despite only having polynomial sample complexity). We give a detailed discussion in Section 5.3.
  • (Notion of a random instance:) A random Hamiltonian phase state |$\ket{\Phi_{\boldsymbol{\theta}}^{\mathbf{A}}}$, e.g., as in Eq. 2.2, is naturally defined in terms of a random binary matrix $\mathbf{A}\leftarrow\mathbb{Z}_{q}^{m\times n}$ and a random set of angles $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{m})$. Hence, it can be efficiently generated by a simple quantum circuit comprising $O(m/n\cdot n^{2})$ CNOT gates, $n$ Hadamard gates, and $m$ single-qubit $Z$ rotations.
  • (Evidence of average-case hardness:) Our learning task admits a worst-case to average-case reduction. In Section 5.1, we separately show how to re-randomize the architecture and the set of angles. Therefore, the hardness of our problem reduces to a worst-case version of the problem.
@@
  • (Cryptographic applications:) Hamiltonian Phase states have a sufficient amount of structure which suffices to construct a number of interesting cryptographic primitives which we sketch in detail in Section 2.3.
@@ 2.3 Applications
The result of [ma2024pseudorandom] show that the ensemble of unitaries, colloquially known as the $\mathsf{PFC}$-ensemble [metger2024simpleconstructionslineardepthtdesigns], form an approximation to a Haar random unitary. However, this construction is not well suited for the HPS assumption, which, in some sense, provides a pseudorandom diagonal unitary. Using the techniques introduced in [ma2024pseudorandom], we show that another pseudorandom construction from [ji2018pseudorandom], the $\mathsf{FHFHFC}$-ensemble is also an approximation to a Haar random unitary when both functions are taken to be uniformly random binary phase oracles. This resolves a conjecture proposed by Ji, Liu and Song [ji2018pseudorandom]. We further show how diagonal matrices can be used to replace the Clifford operation with more applications of a random diagonal matrix and Hadamard gates. With this result, we provide a plausible construction of efficient pseudorandom unitaries from the HPS assumption: alternating applications of HPS unitaries and Hadamards. See Section 6.6 for more details.
@@ Acknowledgements
JB is supported by Henry Yuen’s AFORS (award FA9550-21-1-036) and NSF CAREER (award CCF2144219). JH acknowledges funding from the Harvard Quantum Initiative postdoctoral fellowship. DH acknowledges funding from the US Department of Defense through a QuICS Hartree fellowship. AP is supported by the National Science Foundation (NSF) under Grant No. CCF-1729369.
@@ 6.6 Pseudorandom Unitaries
In this section, we show that a variant of our $\mathsf{HPS}$ assumption allow us construct pseudorandom unitaries. We first show that a unitary that repeatedly applies (uniformly random) phase oracles and Hadamards is indistinguishable from a truly Haar random unitary.
Let us first give a formal definiton of pseudorandom unitaries, as proposed by Ji, Liu, and Song [ji2018pseudorandom]. Broadly speaking, these are efficient ensembles of unitaries which emulate Haar random unitaries to computationally bounded adversaries.

### v3

@@ Abstract
In this work, we seek to make progress on both of these fronts simultaneously—by decoupling quantum pseudorandomness from classical cryptography altogether. We introduce a quantum hardness assumption called the Hamiltonian Phase State ($\mathsf{HPS}$) problem, which is the task of decoding output states of a random instantaneous quantum polynomial-time (IQP) circuit. Hamiltonian phase states can be generated very efficiently using only Hadamard gates, single-qubit $Z$ rotations and CNOT circuits. We show that the hardness of our problem reduces to a worst-case version of the problem, and we provide evidence that our assumption is plausibly fully quantum; meaning, it cannot be used to construct one-way functions. We also show information-theoretic hardness when only few copies of $\mathsf{HPS}$ are available by proving an approximate $t$-design property of our ensemble. Finally, we show that our $\mathsf{HPS}$ assumption and its variants allow us to efficiently construct many pseudorandom primitives, ranging from pseudorandom states, to quantum pseudoentanglement, to pseudorandom unitaries, and even primitives such as public-key encryption with quantum keys.
@@ 2.2 Overview of our Results
  • (Evidence of worst-case hardness) The learnability of ensembles of phase states has been studied extensively in recent work [arunachalam2023optimal], and has been found to have exponential time complexity in the worst case (despite only having polynomial sample complexity). We give a detailed discussion in Section 5.3.
  • (Notion of a random instance) A random Hamiltonian phase state |$\ket{\Phi_{\boldsymbol{\theta}}^{\mathbf{A}}}$, e.g., as in Eq. 2.2, is naturally defined in terms of a random binary matrix $\mathbf{A}\leftarrow\mathbb{Z}_{q}^{m\times n}$ and a random set of angles $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{m})$. Hence, it can be efficiently generated by a simple quantum circuit comprising $O(m/n\cdot n^{2})$ CNOT gates, $n$ Hadamard gates, and $m$ single-qubit $Z$ rotations.
  • (Evidence of average-case hardness) Our learning task admits a worst-case to average-case reduction. In Section 5.1, we separately show how to re-randomize the architecture and the set of angles. Therefore, the hardness of our problem reduces to a worst-case version of the problem.
@@
  • (Cryptographic applications) Hamiltonian Phase states have a sufficient amount of structure which suffices to construct a number of interesting cryptographic primitives which we sketch in detail in Section 2.3.
@@ 2.3 Applications
The result of [ma2024pseudorandom] show that the ensemble of unitaries, colloquially known as the $\mathsf{PFC}$-ensemble [metger2024simpleconstructionslineardepthtdesigns], form an approximation to a Haar random unitary. However, this construction is not well suited for the HPS assumption, which, in some sense, provides a pseudorandom diagonal unitary. [ji2018pseudorandom] presented an alternative construction of pseudorandom unitaries that involves alternating random diagonal gates and $n$-qubit Hadamard gates. We propose that instantiating the random diagonal gates using randomly sampled HPS instances yields a computationally secure pseudo-random unitarySee Section 6.6 for more details.
@@ Acknowledgements
JB is supported by Henry Yuen’s AFORS (award FA9550-21-1-036) and NSF CAREER (award CCF2144219). JH acknowledges funding from the Harvard Quantum Initiative postdoctoral fellowship. DH acknowledges funding from the US Department of Defense through a QuICS Hartree fellowship. AP is supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704.
@@ 6.6 Pseudorandom Unitaries
In this section, we show that a variant of our $\mathsf{HPS}$ assumption might yield pseudorandom unitaries. Let us first give a formal definiton of pseudorandom unitaries, as proposed by Ji, Liu, and Song [ji2018pseudorandom]. Broadly speaking, these are efficient ensembles of unitaries which emulate Haar random unitaries to computationally bounded adversaries.